# Kronecker-product random matrices and a matrix least squares problem

### v2

@@ Appendix A Contour integral representation in the case $\Theta=\Xi=\eta I$
$[(A\otimes I+I\otimes B)^{2}+\eta^{2}I\otimes I]^{-1}=\frac{1}{\eta}\Im[A\otimes I+I\otimes B-i\eta I\otimes I]^{-1},$
@@
[…]
@@ Appendix B Operator concentration inequalities
[…]
@@
By [15, Theorem 1], for a universal constant $C>0$,
[…]
@@
where $j+i-1$ is interpreted modulo $n$. Thus $\hat{\mathsf{x}}$ has $\mathcal{X}$-valued entries $\mathsf{x}_{1},\ldots,\mathsf{x}_{n}$ along the first column, and $T(\hat{\mathsf{x}})$ has these entries instead along the main diagonal. We have
@@ Appendix C Fluctuation averaging
In this section, we prove Lemma 3.5. The proof is analogous to the argument in the scalar setting of [30, Lemma A.2]. For part (c), we will apply the following combinatorial lemma from [30].
@@
By scale invariance of the statement of the lemma, let us assume without loss of generality that $\|\mathbf{u}\|_{\infty}=n^{-1}$. Applying this and the bound |$|\{\mathbf{i}:T(\mathbf{i})=t\}|\leq C_{l}^{\prime}n^{t+(l-t)/2}$ for a constant $C_{l}^{\prime}>0$, we get (for different constants $C_{l},C_{l}^{\prime}>0$)
@@
We collect here several pieces of background on von Neumann algebras and non-commutative $L^{p}$ spaces that are needed in our main arguments. We refer to [50], [14, Section 1], and [58, Chapter 14] for additional discussion. Throughout, $\mathcal{X}$ is a (finite) von Neumann algebra with faithful, normal, tracial state $\phi$ as in Section 3.
@@
Defining $\|\mathsf{x}\|_{p}=\phi(|x|^{p})^{1/p}$, the space $L^{p}(\mathcal{X})$ is the Banach space completion of $\mathcal{X}$ under $\|\cdot\|_{p}$. We set $\|\mathsf{x}\|_{\infty}\equiv\|\mathsf{x}\|_{\mathrm{op}}$ and $L^{\infty}(\mathcal{X})\equiv\mathcal{X}$. These spaces $L^{p}(\mathcal{X})$ may be continuously embedded into a common space of (unbounded) densely-defined operators affiliated to $\mathcal{X}$ — we refer to [50] or [14, Section 1] for this construction.
@@
$\|\mathsf{x}\|_{p}\leq\|\mathsf{y}\|_{p}\text{ for all }\mathsf{x},\mathsf{y}\in L^{p}(\mathcal{X})\text{ with }0\leq\mathsf{x}\leq\mathsf{y},\qquad|\phi(x)|\leq\|\mathsf{x}\|_{1}\text{ for all }\mathsf{x}\in L^{1}(\mathcal{X}).$

### v3

@@ Appendix A Contour integral representation in the case $\Theta=\Xi=\eta I$
$(A\otimes I+I\otimes B)^{2}+\eta^{2}I\otimes I]^{-1}=\frac{1}{\eta}\Im[A\otimes I+I\otimes B-i\eta I\otimes I]^{-1},$
@@
$=\frac{1}{2\pi i}\oint_{\Gamma}\frac{1}{w-\lambda_{j}}\frac{1}{w+\mu_{k}-z}\mathbf{u}_{j}\mathbf{u}_{j}^{*}\otimes\mathbf{v}_{k}\mathbf{v}_{k}^{*}\,dw$
$={-}\frac{1}{2\pi i}\oint_{\Gamma}R_{A}(w)\otimes R_{B}(z-w)\,dw$
@@ Appendix B Operator concentration inequalities
Let $(\alpha_{i})_{i=1}^{n}$ be a sequence of independent $\mathbb{C}$-valued random variables, let $(\alpha_{i}^{\prime})_{i=1}^{n}$ be an independent copy of $(\alpha_{i})_{i=1}^{n}$, and let $(\mathsf{y}_{ij}:i,j=1,\ldots,n)$ be elements of a Banach space with norm $\|\cdot\|$. Then for a universal constant $C>0$,
@@
By [16, Theorem 1], for a universal constant $C>0$,
$\frac{1}{p}\,\mathbb{E}\bigg[\bigg\|\sum_{i\neq j}\alpha_{i}\alpha_{j}\mathsf{y}_{ij}\bigg\|^{p}\bigg]$ $=\int_{0}^{\infty}t^{p-1}\mathbb{P}\bigg[\bigg\|\sum_{i\neq j}\alpha_{i}\alpha_{j}\mathsf{y}_{ij}\bigg\|\geq t\bigg]\;dt$
$\leq\int_{0}^{\infty}Ct^{p-1}\mathbb{P}\bigg[C\bigg\|\sum_{i\neq j}\alpha_{i}\alpha_{i}^{\prime}\mathsf{y}_{ij}\bigg\|\geq t\bigg]\;dt=\frac{C^{p+1}}{p}\mathbb{E}\bigg[\bigg\|\sum_{i\neq j}\alpha_{i}\alpha_{i}^{\prime}\mathsf{y}_{ij}\bigg\|^{p}\bigg].$
@@
[…]
@@ Appendix C Fluctuation averaging
In this section, we prove Lemma 3.5. The proof is analogous to the argument in the scalar setting of [29, Lemma A.2]. For part (c), we will apply the following combinatorial lemma from [29].
@@
By scale invariance of the statement of the lemma, let us assume without loss of generality that $\|\mathbf{u}\|_{\infty}=n^{-1}$. Applying this and the bound $|\{\mathbf{i}:T(\mathbf{i})=t\}|\leq C_{l}^{\prime}n^{t+(l-t)/2}$ for a constant $C_{l}^{\prime}>0$, we get (for different constants $C_{l},C_{l}^{\prime}>0$)
@@
We collect here several pieces of background on von Neumann algebras and non-commutative $L^{p}$ spaces that are needed in our main arguments. We refer to [50], [15, Section 1], and [58, Chapter 14] for additional discussion. Throughout, $\mathcal{X}$ is a (finite) von Neumann algebra with faithful, normal, tracial state $\phi$ as in Section 3.
@@
Defining $\|\mathsf{x}\|_{p}=\phi(|x|^{p})^{1/p}$, the space $L^{p}(\mathcal{X})$ is the Banach space completion of $\mathcal{X}$ under $\|\cdot\|_{p}$. We set $\|\mathsf{x}\|_{\infty}\equiv\|\mathsf{x}\|_{\mathrm{op}}$ and $L^{\infty}(\mathcal{X})\equiv\mathcal{X}$. These spaces $L^{p}(\mathcal{X})$ may be continuously embedded into a common space of (unbounded) densely-defined operators affiliated to $\mathcal{X}$ — we refer to [50] or [15, Section 1] for this construction.
@@
[…]